Question B onlyDavis Kitchen Supply produces stoves for commercial kitchens. The costs to manufacture and market the stoves at the companysa. Market research estimates that volume could be increased to 7,000 units, which is well within production capacity limitatiWithout Government Contract Impact With Government Contract Regular Government Total Revenue Variable manufacturing costs Var
Answer #1


Davis Kitchen Supply Requirement al Impact Sales Price Quantity Revenue Variable manufacturing costs Variable marketing costs

А 3 Requirement a1 Impact Before Price Reduction 386 6000 =B6*B7 =174*B7 =33*B7 =B8-B9-B10 =68*B7 =78*B7 =B11-B12-B13 6 Sales